Suppose that a teacher driving a 1972 LeMans zooms out of a darkened tunnel at 29.3 m/s. He is momentarily blinded by the sunshine. When he recovers, he sees that he is fast overtaking a bus ahead in his lane moving at the slower speed of 14.3 m/s. He hits the brakes as fast as he can (his reaction time is 0.31 s). If he can decelerate at 3.3 m/s2, what is the minimum distance between the driver and the bus when he first sees it so that they do not collide?
1
Expert's answer
2020-09-15T10:04:03-0400

Given,

Speed of the car "(u)=29.3 m\/sec"

Speed of bus "(U)=14.3 m\/sec"

Reaction time "(t_r)=0.31 sec"

Deceleration "(a)=-3.3 m\/sec^2"

Let the time gap between the bus and car after seeing the bus is t,

Let the distance covered by bus in t time is "x" and initial distance between bus and car is "D"

So, "x= U\times t = 14.3 t ------(i)"

distance covered by car in the same duration of time "D+x=u\times t+(u(t-0.31)-\dfrac{a(t-0.31)^2}{2}) -----(ii)"

From equation (i) and (ii),

Substituting the value of "x," in equation (ii)

"D+14.3t=u\times t+(u(t-0.31)-\dfrac{a(t-0.31)^2}{2})"

"\Rightarrow D=29.3\times 0.31 -14.3t+29.3t-29.3\times 0.31-\dfrac{3.3(t-0.31)^2}{2}"

"\Rightarrow D=15t-\dfrac{3.3(t-0.31)^2}{2}"

taking the differentiation of both side,

"\dfrac{dD}{dt}=15-3.3(t-0.31)"

for maxima and minima,

"\dfrac{dD}{dt}=0"

"\Rightarrow 15-3.3(t-0.31)=0"

"\Rightarrow 3.3(t-0.31)=15"

"\Rightarrow t-0.31=\dfrac{15}{3.3}"

"\Rightarrow t=(0.31+4.54) sec"

"\Rightarrow t =4.85 sec"

Hence, minimum distance between the car and bus

"\Rightarrow D=15\times 4.85- \dfrac{3.3(4.85-0.31)^2}{2}"

"\Rightarrow D =(72.75-33.41)m"

"\Rightarrow D=39.34 m"